# Effort-aware Fairness: Incorporating a Philosophy-informed, Human-centered Notion of Effort into Algorithmic Fairness Metrics

### v1

@@ Introduction
Figure 3: Effort-aware Group Parity (in terms of mean predicted risk) as a function of Effort (SHED data). Bin length is 0.1. Only demographic groups with at least 10 data points are considered. Random forest consistently achieves the best EaGF results in terms of Race and Age, but not Sex.
Figure 3 shows effort-aware, i.e., within-bin, demographic group parity as a function of effort. To illustrate why EaGF information is helpful, we first look at Table 6 (Appendix F) on overall demographic parity, i.e., not conditioned on similar efforts, showing random forest to be among the most group-wise fair models in terms of race, sex, or age group. Figures 3(a) and 3(c) reassure us that conditioned on similar efforts, random forest remains the fairest model among racial and age groups. However, Figure 3(b) cautions us that when conditioned on higher-effort bins ($0.4$-$0.5$ and $0.5$-$0.6$), random forest is among the least fair models in terms of sex. This pattern might be concerning since we do not want to disincentivize people from exerting more effort. If sex parity is a legal/policy priority in our context, our EaGF audit might suggest using XGBoost instead of random forest because XGBoost achieves roughly the same best overall sex parity as random forest (Table 6) but XGBoost achieves the best EaGF score when conditioned on the highest-effort bin (Figure 3(b)), thereby incentivizing people to exert more efforts. Our demonstration illustrates a key strength of EaGF compared to traditional group fairness as this metric allows a more granular view on whether there is any concerning trend of parity vs. effort when deciding among many AI models.
@@ Conclusion
We identify a gap between how the CS/AI literature has attempted to formalize the notion of “effort” into existing fairness metrics and how philosophers/legal scholars understand “efforts”. We seek to fill this gap by operationalizing the Force-based account of Effort from philosophy, formulating an “effort” metric as the product of inertia (systemic disadvantage beyond an individual’s control, such as childhood poverty) and accelerations (temporal changes in a task-related input-space feature), which can be looped into existing fairness metrics to formulate Effort-aware Group/Individual Fairness (EaGF/EaIF). As accelerations can be conceptualized more broadly as trajectory, we conduct a human subjects experiment and successfully validate that the acceleration/trajectory component of our Effort-aware Fairness formulation aligns with laypeople’s perception of fairness across two distinct AI-assisted decision-making contexts, namely criminal justice and personal finance. Finally, we develop practical pipelines to compute EaIF and EaGF on real-world datasets (SHED and CLUE), demonstrating how these metrics complement traditional fairness metrics.
Ethical considerations statement
@@
We minimize privacy risks to our human subjects experiment participants by removing any identifiable information (such as Prolific ID) from any supplementary materials we publish, as well as complying with any related institutional (IRB) and federal human subjects protection policies. We commit to compensating our participants reasonably. For example, as our original estimation of 30 minutes for study completion time turns out to be an under-estimation compared to the median completion time (42 minutes), we have adjusted the pay rate from \$$\$6$ to \$$\$8.45$ to maintain a median hourly compensation rate of \$$\$12$, which is considered a “Good” rate in Prolific standard. We also carefully comply with the CLUE dataset’s term of use to protect defendants’ privacy, such as by storing the dataset on a password-secured server and only using non-personally identifiable features in our simple AI models development and fairness evaluation.
@@ Appendix A Appendix A: Judicial Interpretations of Equal Protection: Bridging Group and Individual Fairness
Case law deems both group fairness and individual fairness as relevant to the constitutional “Equal Protection” clause. For example, the U.S. Supreme Court (SCOTUS) rules in Washington v. Davis (1976) that: “Disproportionate impact is not irrelevant, but it is not the sole touchstone of an invidious racial discrimination forbidden by the Constitution.” (SCOTUS 1976). Older SCOTUS decisions such as F.S. Royster Guano Co. v. Commonwealth of Virginia (1920) interpret “Equal protection” consistently with individual fairness: “The Equal Protection Clause of the Fourteenth Amendment commands […] that all persons similarly situated should be treated alike” (SCOTUS 1920). Interestingly, a synergy of both group and individual fairness is found in City of Cleburne, Tex. v. Cleburne Living Center (1985): “Discrimination, in the Fourteenth Amendment sense, connotes a substantive constitutional judgment that two individuals or groups are entitled to be treated equally with respect to something” (SCOTUS 1985).
@@ Appendix B Appendix B: Discussion about Counterfactual Effort in AI Fairness Literature
Our main concern with this AI Fairness literature is that “effort” was framed but should not have been framed in a counterfactual sense as recourse or the distance (in the input feature space) between a decision subject’s current position and the decision boundary (which separates feature vectors that get good outcomes from feature vectors that get bad outcomes). We recognize the importance of these counterfactual “effort” fairness or recourse fairness metrics, such as “equal improvability” (Guldogan et al. 2023), which often seek to equalize the distance to the decision boundary across demographic groups.
However, defining effort in terms of recourse has two theoretical shortcomings. First, recourse might vary depending on the AI model, as the decision boundary will change with the model. However, effort should be model-agnostic and should depend only on the personal history of each individual. Second, “effort” as recourse cannot be incorporated into the individual fairness regime established by Dwork et al. (2012) because recourse is only well-defined for decision subjects receiving bad outcomes, not for decision subjects already receiving good outcomes. Suppose we incorporate “effort” as recourse into the individual distance function (Step 1 of individual fairness evaluation). If that distance function identifies two individuals with highly similar recourse, those recourse values must be well-defined and the outcomes of those two individuals must be the same (both good), automatically guaranteeing a perfect individual fairness score. We give an example to illustrate the incompatibility between recourse and the individual fairness regime: Von Kügelgen et al. (2022) defines an “Individually Fair Causal Recourse” metric as “if the cost of recourse would have been the same had the individual belonged to a different protected group.” However, there is no individual distance function defined, and this metric is more rooted in procedural fairness (a protected demographic feature should not impact a quantity of interest, such as cost of recourse) than in “individual fairness” as defined by Dwork et al. (2012).
@@ Appendix C Appendix C: Social Science Grounding of Effort-aware Fairness
More fundamentally, we acknowledge an inherent limitation in using a concept rooted in the natural sciences (force) to characterize a social phenomenon (human effort). Prior AI fairness literature also shows little of such intellectual borrowing. As much as we believe that multidimensional social phenomena can never have a straightforward quantification, we also believe in the power of analogical reasoning, and in this particular case, we believe our work’s analogical transfer has great potential to be fruitful. Though we are the first to operationalize effort as force in the context of algorithmic fairness, we are not the first to lean on the powers of analogical reasoning to facilitate social science theory development, e.g.,“Bytes provide a basis for grounding the study of social relationship and institutions in the natural sciences because their transmission or storage requires changes in energy or physical states” (Turnbull 2002). We do not claim to provide an exact translation of effort into force; rather, we aim for this characterization to be a productive metaphor that can facilitate new debate in the community on how to better measure effort in the AI fairness context.
@@ Appendix D Appendix D: Details on Predictive Models
For CLUE predictive models, we build predictive features using three years of history (which spans data from $2013$ to $2018$), including current case details, pending cases, past arrests and convictions across degrees (e.g., misdemeanor, felony), categories (e.g., DUI, drug, public order, traffic criminal, violent), and in specific time frames (e.g., past $30$ days, …, $1$ year, $4$ years). We set the high recidivism risk threshold to be the $62^{th}$ percentile of predicted risk scores across the felony offenders of the $25,000$ defendants in our cleaned CLUE version. This choice is inspired by the finding from the “Pretrial Release of Felony Defendants in State Courts” report by the DOJ (2007) [^12] https://bjs.ojp.gov/content/pub/pdf/prfdsc.pdf that $62\%$ of felony defendants between $1990-2004$ were released before trial, which might imply that roughly $38\%$ of felony defendants in that period were considered high risk. The “high risk” threshold we get with this method is roughly $0.40$.
@@
For the SHED outcome variable, the respondent’s frequency of unpaid credit card balance is recorded as the answer to the question in the survey: “In the past 12 months, how frequently have you carried an unpaid balance on one or more of your credit cards?” The responses are encoded as values from $\{0,1,2,3\}$ where 0 indicates “Never carried an unpaid balance (always pay in full),” $1$ represents “Once,” $2$ corresponds to “Some of the time,” and 3 signifies “Most or all of the time.” This feature is then converted into a binary variable, where 00 represents “Never carried an unpaid balance (always pay in full)” and $1$ indicates “Miss credit card balance payment at least once.”
@@
The features considered include income data from $2019$ to $2022$, employment status (e.g., employee, self-employed, full-time, part-time), household composition (e.g., the number of individuals living with the household head), demographics (e.g., education level, age), and financial status (e.g., whether the household has a checking, savings, or money market account). We split the data into a training set ($70\%$) and a testing set ($30\%$). Based on the evaluation metrics, Logistic Regression outperforms XGBoost and Random Forest in predicting missed payments. Logistic regression achieves an accuracy of 72.46% with a weighted F1 score of 0.71. In comparison, the best-performing XGBoost model—tuned via randomized search—achieves an accuracy of 64.25% and a weighted F1 score of 0.64, while the random forest model achieves an accuracy of 65.70% with a weighted F1 score of 0.64. To obtain out-of-sample risk scores for all households, we use 5-fold cross-validation. This involves dividing the dataset into $5$ disjoint subsets. In each iteration, one subset is treated as test data, while the remaining $4$ subsets are used to train the model and generate predictions for the test subset. This process is repeated $5$ times, ensuring predictions are obtained for all subsets.
@@ Appendix E Appendix E: Human Subjects Experiment Interfaces and Incentives
Figure 4: Example human subjects experiment interfaces to measure: 1/ trajectory distance (upper left); 2/ aggregate distance (upper right); 3/ overall input-space distance (lower left); 4/ output-space distance (lower right). Pair: ($K_{1},D_{1}$). Condition: CLUE.
To incentivize participants to give a truthful answer to a multiple-choice question (MCQ), this method adds a follow-up question, asking each participant to predict the frequency of every answer choice in the main MCQ as selected by other participants. For each pair of MCQ and follow-up prediction question, each participant gets a BTS score = information score + prediction score. The first term, information score, will be high if the participant chooses a “surprisingly common” answer in the MCQ, i.e., the predicted frequency of the participant’s answer (averaged across all participants’ predictions) is higher than that answer’s actual frequency. The intuition behind this information score is that people tend to overestimate the frequency of an answer if they think the answer is the truth. The second term, prediction score, simply rewards each individual participant’s prediction if it is accurate, i.e., close to the actual frequency. The BTS method often informs participants, without revealing the math to them, that they will receive bonuses if they provide the most “truthful” answers, which, behind the scene, means if they get the highest BTS scores across their answers. We give an example BTS question in Figure 5.
@@ Appendix F Appendix F: Supporting Figures and Tables
$\Delta r_{\text{input}}$=$r_{\text{traj},\text{input}}-r_{\text{aggr},\text{input}}$ $[-0.2780,-0.1280]$ $[-0.1950,-0.0730]$
$\Delta r_{\text{output}}$=$r_{\text{traj},\text{output}}-r_{\text{aggr},\text{output}}$ $[-0.3000,-0.1280]$ $[-0.3330,-0.1610]$
@@
$d_{E}\times C_{\text{CLUE}}$0.0661*** -0.0163***
@@ Appendix G Appendix G: Demonstrating Effort-aware Fairness metrics on CLUE dataset
Regarding EaGF, although Table 8 on traditional, effort-unaware demographic parity may give an impression that decision tree achieves the best group fairness across all three demographic features (race, sex, and age group), more nuanced investigation of EaGF in Figure 6 shows that decision tree only achieves consistently the best EaGF results among age groups. Between sexes, while decision tree achieves the best EaGF at the lower-effort bins (00 to $0.3$), Light GBM achieves the best EaGF at most of the higher-effort bins ($0.3$ to $0.9$), indicating that Light GBM might be a better fit if the application prioritizes equality between males and females of similar efforts while incentivizing both sexes to exert more future efforts. With respect to race, all three models seem to give the right incentive where higher-efforts bins generally suffer less demographic disparity.
@@
Figure 6: Effort-aware Group Parity (in terms of mean predicted risk) as a function of Effort (CLUE data). Bin length is 0.1. Only demographic groups with at least 10 data points are considered. Decision tree achieves the best age group parity (compared to logistic regression and light GBM) in most effort bins. Higher-effort bins get better race parity across all three AI models.

### v3

@@ Introduction
Figure 3: Effort-aware Group Parity (in terms of mean predicted risk) as a function of Effort (SHED data). Bin length is 0.1. Only demographic groups with at least 10 data points are considered. The subplots’ x-axes might differ since not all Effort bins within the 0-1 range have enough data points to compute EaGF. Random forest consistently achieves the best EaGF results in terms of race and age, but not sex.
Figure 3 shows Effort-aware, i.e., within-bin, demographic group parity as a function of Effort. To illustrate why EaGF information is helpful, we first look at Table 6 (Appendix F) on overall demographic parity, i.e., not conditioned on similar Efforts, showing random forest to be among the most group-wise fair models in terms of race, sex, or age group. Figures 3(a) and 3(c) reassure us that, conditioned on similar Efforts, random forest remains the fairest model among racial and age groups. However, Figure 3(b) cautions us that when conditioned on higher-effort bins (0.40.4-$0.5$ and 0.50.5-$0.6$), random forest is among the least fair models in terms of sex. This pattern might be concerning since we do not want to disincentivize people from exerting more Effort. If sex parity is a legal/policy priority in our context, our EaGF audit might suggest using XGBoost instead of random forest because XGBoost achieves roughly the same best overall sex parity as random forest (Table 6) and XGBoost achieves the best EaGF score when conditioned on the highest-Effort bin (Figure 3(b)), thereby incentivizing people to exert more effort. Our demonstration highlights a key strength of EaGF over traditional group fairness, as it provides a more granular view of any concerning trend in parity vs. Effort when evaluating AI models.
@@ Conclusion
We identify a gap between how the CS/AI literature has attempted to formalize the notion of “Effort” into existing fairness metrics and how philosophers/legal scholars understand “Efforts”. We seek to fill this gap by operationalizing the Force-based account of Effort from philosophy, formulating an “Effort” metric as the product of inertia (systemic disadvantage beyond an individual’s control, such as childhood poverty) and accelerations (temporal changes in a task-related input-space feature), which can be looped into existing fairness metrics to formulate Effort-aware Group/Individual Fairness (EaGF/EaIF). As we can conceptualize acceleration more broadly as trajectory, we conduct a human subjects experiment and successfully validate that the acceleration/trajectory component of our Effort-aware Fairness formulation aligns with laypeople’s perception of fairness across two distinct AI-assisted decision-making contexts, namely criminal justice and personal finance. Finally, we develop practical pipelines to compute EaIF and EaGF on real-world datasets (SHED and CLUE), demonstrating how these metrics complement traditional fairness metrics.
Ethics Statement
@@
We minimize privacy risks to our human subjects experiment participants by removing any identifiable information (such as Prolific ID) from any supplementary materials we publish, as well as complying with any related institutional (IRB) and federal human subjects protection policies. We commit to compensating our participants reasonably. For example, as our original estimation of 30 minutes for study completion time turns out to be an under-estimation compared to the median completion time (42 minutes), we have adjusted the pay rate from \$6\\$6 to \$8.45\\$8.45 to maintain a median hourly compensation rate of \$12\\$12, which is considered a “Good” rate in Prolific standard. We also carefully comply with the CLUE dataset’s terms of use to protect defendants’ privacy, such as by storing the dataset on a password-secured server and only using non-personally identifiable features in our simple AI models development and fairness evaluation.
We acknowledge two core limitations of our proposed metrics. First, in our formulation of both EaIF and EaGF, by using efforts to find pairs of similar individuals or by conditioning on similar-effort bins when computing group parity metrics, we normatively prescribe the ideal outcomes of similar-effort individuals. Still, we leave open the more complex normative question of how the outcomes of individuals with different levels of Effort should be. We have not incorporated a notion of directionality for the relationship between Effort and outcome. In other words, we do not necessarily stipulate that people who exert more Effort should always receive more favorable outcomes, as this involves more unsettled philosophical debates on meritocracy, e.g., whether resources should be allocated based on one’s efforts or one’s natural talents (such as a basketball athlete’s height), which are often due to luck and beyond one’s control (Sandel 2021). Second, the risk of fairwashing, i.e., generating posthoc explanations to make predictions by an otherwise unfair black-box AI models score well on specific fairness metrics (A¨ıvodji et al. 2019, 2021), might extend to our EaF metrics because bad actors might manipulate the particular time steps when computing our acceleration term as well as the feature(s) from which our inertia term or the societal holding back effect is approximated. Therefore, we welcome future research on how to guard the quantification of Efforts against fairwashing, as well as more AI-assisted decision-making data collection initiatives to include more individual-level features that might reflect disadvantages beyond one’s control, such as disability or childhood poverty.
Acknowledgments
This material is based on work supported in part by the Institute for Trustworthy AI in Law and Society (TRAILS), which is supported by the National Science Foundation under Award No. 2229885. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or the National Institute of Standards and Technology. This work is also supported in part by the National Institute of Justice’s Graduate Research Fellowship (Award No. 15PNIJ-23-GG-01932-RESS).
@@ Appendix A Appendix A: Judicial Interpretations of Equal Protection: Bridging Group and Individual Fairness
Case law deems both group fairness and individual fairness as relevant to the constitutional “Equal Protection” clause. For example, the U.S. Supreme Court (SCOTUS) ruled in Washington v. Davis (1976) that: “Disproportionate impact is not irrelevant, but it is not the sole touchstone of an invidious racial discrimination forbidden by the Constitution.” (SCOTUS 1976). Older SCOTUS decisions such as F.S. Royster Guano Co. v. Commonwealth of Virginia (1920) interpret “Equal protection” consistently with individual fairness: “The Equal Protection Clause of the Fourteenth Amendment commands […] that all persons similarly situated should be treated alike” (SCOTUS 1920). Interestingly, a synergy of both group and individual fairness is found in City of Cleburne, Tex. v. Cleburne Living Center (1985): “Discrimination, in the Fourteenth Amendment sense, connotes a substantive constitutional judgment that two individuals or groups are entitled to be treated equally with respect to something” (SCOTUS 1985).
@@ Appendix B Appendix B: Discussion about Counterfactual Effort in AI Fairness Literature
Our primary concern with this AI Fairness literature is that “Effort” was framed but should not have been framed in a counterfactual sense as recourse or the distance (in the input feature space) between a decision subject’s current position and the decision boundary (which separates feature vectors that get good outcomes from feature vectors that get bad outcomes). We recognize the importance of these counterfactual “Effort” fairness or recourse fairness metrics, such as “equal improvability” (Guldogan et al. 2023), which often seek to equalize the distance to the decision boundary across demographic groups.
However, defining Effort in terms of recourse has two theoretical shortcomings. First, recourse might vary depending on the AI model, as the decision boundary will change with the model. However, Effort should be model-agnostic and should depend only on the personal history of each individual. Second, “Effort” as recourse cannot be incorporated into the individual fairness regime established by Dwork et al. (2012) because recourse is only well-defined for decision subjects receiving bad outcomes, not for decision subjects already receiving good outcomes.
Suppose we incorporate “Effort” as recourse into the individual distance function (Step 1 of individual fairness evaluation). If that distance function identifies two individuals with highly similar recourse, those recourse values must be well-defined and the outcomes of those two individuals must be the same (both good), automatically guaranteeing a perfect individual fairness score. We give an example to illustrate the incompatibility between recourse and the individual fairness regime: Von Kügelgen et al. (2022) defines an “Individually Fair Causal Recourse” metric as “if the cost of recourse would have been the same had the individual belonged to a different protected group.” However, there is no individual distance function defined, and this metric is more rooted in procedural fairness (a protected demographic feature should not impact a quantity of interest, such as cost of recourse) than in “individual fairness” as defined by Dwork et al. (2012).
@@ Appendix C Appendix C: Social Science Grounding of Effort-aware Fairness
More fundamentally, we acknowledge an inherent limitation in using a concept rooted in the natural sciences (force) to characterize a social phenomenon (human Effort). Prior AI fairness literature also shows little of such intellectual borrowing. As much as we believe that multidimensional social phenomena can never have a straightforward quantification, we also believe in the power of analogical reasoning, and in this particular case, our work’s analogical transfer has great potential to be fruitful. Though we are the first to operationalize Effort as force in the context of algorithmic fairness, we are not the first to lean on the powers of analogical reasoning to facilitate social science theory development, e.g., “Bytes provide a basis for grounding the study of social relationship and institutions in the natural sciences because their transmission or storage requires changes in energy or physical states” (Turnbull 2002). We do not claim to provide an exact translation of Effort into force; instead, we aim for this characterization to be a productive metaphor that can facilitate new debate in the community on how to measure Effort in the AI fairness context better.
@@ Appendix D Appendix D: Details on Predictive Models
For CLUE predictive models, we build predictive features using three years of history (which spans data from $2013$ to $2018$), including current case details, pending cases, past arrests and convictions across degrees (e.g., misdemeanor, felony), categories (e.g., DUI, drug, public order, traffic criminal, violent), and in specific time frames (e.g., past $30$ days, …, $1$ year, $4$ years). We set the high recidivism risk threshold to be the $62^{th}$ percentile of predicted risk scores across the felony offenders of the $25,000$ defendants in our cleaned CLUE version. This choice is inspired by the finding from the “Pretrial Release of Felony Defendants in State Courts” report by the DOJ (2007) [^19] https://bjs.ojp.gov/content/pub/pdf/prfdsc.pdf that $62\%$ of felony defendants between 199020041990-2004 were released before trial, which might imply that roughly $38\%$ of felony defendants in that period were considered high risk. The “high risk” threshold we get with this method is approximately $0.40$.
@@
For the SHED outcome variable, the respondent’s frequency of unpaid credit card balance is recorded as the answer to the question in the survey: “In the past 12 months, how frequently have you carried an unpaid balance on one or more of your credit cards?” The responses are encoded as values from $\{0,1,2,3\}$ where 0 indicates “Never carried an unpaid balance (always pay in full),” $1$ represents “Once,” $2$ corresponds to “Some of the time,” and 3 signifies “Most or all of the time.” This feature is then converted into a binary variable, where 0 represents “Never carried an unpaid balance (always pay in full)” and $1$ indicates “Miss credit card balance payment at least once.”
@@
The features considered include income data from $2019$ to $2022$, employment status (e.g., employee, self-employed, full-time, part-time), household composition (e.g., the number of individuals living with the household head), demographics (e.g., education level, age), and financial status (e.g., whether the household has a checking, savings, or money market account). We split the data into a training set ($70\%$) and a testing set ($30\%$). Based on the evaluation metrics, Logistic Regression outperforms XGBoost and Random Forest in predicting missed payments. Logistic regression achieves an accuracy of 72.46% with a weighted F1 score of 0.71. In comparison, the best-performing XGBoost model—tuned via randomized search—achieves an accuracy of 64.25% and a weighted F1 score of 0.64. In contrast, the random forest model achieves an accuracy of 65.70% with a weighted F1 score of 0.64. To obtain out-of-sample risk scores for all households, we use 5-fold cross-validation. This validation method involves dividing the dataset into $5$ disjoint subsets. In each iteration, one subset is treated as test data, while the remaining $4$ subsets are used to train the model and generate predictions for the test subset. This process is repeated $5$ times, ensuring predictions are obtained for all subsets.
@@ Appendix E Appendix E: Human Subjects Experiment Interfaces and Incentives
Figure 4: Example human subjects experiment interfaces to measure: (1) trajectory distance (upper left); (2) aggregate distance (upper right); (3) overall input-space distance (lower left); (4) output-space distance (lower right). Pair: ($K_{1},D_{1}$). Condition: CLUE
To incentivize participants to give a truthful answer to a multiple-choice question (MCQ), this method adds a follow-up question, asking each participant to predict the frequency of every answer choice in the main MCQ as selected by other participants. For each pair of MCQ and follow-up prediction question, each participant gets a BTS score = information score + prediction score. The first term, information score, will be high if the participant chooses a “surprisingly common” answer in the MCQ, i.e., the predicted frequency of the participant’s answer (averaged across all participants’ predictions) is higher than that answer’s actual frequency. The intuition behind this information score is that people tend to overestimate the frequency of an answer if they think the answer is the truth. The second term, prediction score, rewards each participant’s prediction if it is accurate, i.e., close to the actual frequency. The BTS method often informs participants, without revealing the math to them, that they will receive bonuses if they provide the most “truthful” answers, which, behind the scenes, means if they get the highest BTS scores across their answers. We give an example BTS question in Figure 5.
@@ Appendix F Appendix F: Supporting Figures and Tables
$\Delta r_{\text{input}}$$r_{\text{traj},\text{input}}-r_{\text{aggr},\text{input}}$ [0.2780,0.1280][-0.2780,-0.1280] [0.1950,0.0730][-0.1950,-0.0730]
$\Delta r_{\text{output}}$$r_{\text{traj},\text{output}}-r_{\text{aggr},\text{output}}$ [0.3000,0.1280][-0.3000,-0.1280] [0.3330,0.1610][-0.3330,-0.1610]
@@
$d_{E}\times C_{\text{CLUE}}$ -0.0661*** -0.0163***
@@ Appendix G Appendix G: Demonstrating Effort-aware Fairness metrics on CLUE dataset
Regarding EaGF, although Table 8 on traditional, Effort-unaware demographic parity may give an impression that decision tree achieves the best group fairness across all three demographic features (race, sex, and age group), more nuanced investigation of EaGF in Figure 6 shows that decision tree only achieves consistently the best EaGF results among age groups. Between sexes, while decision tree achieves the best EaGF at the lower-Effort bins (0 to $0.3$), Light GBM achieves the best EaGF at most of the higher-Effort bins ($0.3$ to $0.9$), indicating that Light GBM might be a better fit if the application prioritizes equality between males and females of similar Efforts while incentivizing both sexes to exert more future Efforts. With respect to race, all three models seem to give the right incentive, where higher-effort bins generally suffer less demographic disparity.
@@
Figure 6: Effort-aware Group Parity (in terms of mean predicted risk) as a function of Effort (CLUE data). Bin length is 0.1. Only demographic groups with at least 10 data points are considered. The subplots’ x-axes might differ since not all Effort bins within the 0-1 range have enough data points to compute EaGF. Decision tree achieves the best age group parity (compared to logistic regression and light GBM) in most Effort bins. Higher-Effort bins get better race parity across all three AI models.